Suppose there are just two countries, A and B, in the oil market and the inverse demand for oil is given by P = 100 – Q. Both countries have the same marginal cost of producing oil which is €40.

The price that each country would charge in the Cournot equilibrium is

€40

€70

€60

€30
Answer #1

Answer:

A. First of all we need to find out Marginal Revenue equation from total revenue equation

Total revenue = P * Q

= ( 100 - Q ) Q

= 100Q - Q^2

Thus total revenue = 100Q - Q^2

Marginal revenue by derivation = 100 - 2Q from total revenue

B. Then we need to find optimum quantity

Here MC = 40

For profit maximizing MR = MC

Thus, 100 - 2Q = 40

2Q = 100 - 40

Q = 60/2 = 30

Thus optimum quantity is 30 units

Now we have quantity, we can find out price

Price = 100 - Q

= 100 - 30 = €70

Thus €70 price will be charged